OpenStudy (anonymous):

what are the approximate solutions of 2x^2-x+10=0? f) -2,2.5 g) -1.97,2.47 h)-2.5,2 i) no solution. help me please. :*

OpenStudy (anonymous):

do you know the quadratic formula?

OpenStudy (anonymous):

yeah but i don't know how to use it on a question like this. :(

OpenStudy (anonymous):

ok ... let's give it a try ;) \[\LARGE {x_{1/2}} = \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}}\] \[\LARGE \Delta=b^2-4ac\] if \[\LARGE \Delta>0 \;\;two \;\;solutions\] \[\LARGE \Delta=0 \;\;one \;\;solution\] \[\LARGE \Delta<0 \;\;no\;\;solution \]

OpenStudy (anonymous):

now... applying the given ... we get: \[\LARGE {x_{1/2}} = \frac{{ - 1 \pm \sqrt {1 - 4 \cdot 20} }}{{2a}}\] from here we see that expression under square root will be negative... and therefore it has no solution... :)

OpenStudy (anonymous):

so the answer is i) no solution

OpenStudy (anonymous):

omg, thank youuuu. i'll have other questions up about this cause i'm like totally lost, but thank you so much. <3

OpenStudy (anonymous):

where are you stucked? ... post it on wall. (not here) so other members can help you, (If I don't know it) .. :)